# Part I: Nonlinear Codes

### v1

@@ V Joining Cycles in the Factor Graph
Our next step in combining the codewords is to order the $2^{k(2^{t}-1)}$ elements of ${\cal Q}_{1}$ in a cyclic Gray code, i.e., that each two elements will differ in exactly one position. Such a cyclic Gray code is well-known [38] and is also very easy to construct. The following lemma can be readily verified.
@@
can be combined into one codeword by the merge-or-split method, i.e., they have a pair of companion vertices.
@@
The two codewords in $Q_{1}$ differ in exactly one coordinate implies that the associated two codewords of $\mathbb{C}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, differ in exactly two coordinates separated by $12^{t}-$ 1 coordinates in which both ${\cal C}_{1}$ and ${\cal C}_{2}$ are identical. The two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ associated with these $12^{t}-$ 1 coordinates and the coordinate which follows them differ only in the last column. This implies that these two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ are companion vertices. ∎
The cyclic Gray code of the elements of the set ${\cal Q}_{1}$ and Lemma 14 implies that we can form a cyclic Gray code from the codewords of each equivalence class.
The next step is to order the codewords from different equivalence classes associated with the same self-dual sequence ${\cal S}$. For this purpose we define another Gray code on the set of elements $Q_{2}$ defined by
@@
It is easy to verify again where we can find a pair of companion vertices in two distinct equivalence classes associated with the same self-dual sequence ${\cal S}$, which differ in exactly one coordinate of $Q_{2}$.
@@
If $(i_{i},i_{2},i_{3},\ldots,i_{2^{t}})\in{\cal Q}_{2}$ and $(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3},\ldots,i^{\prime}_{2^{t}})\in{\cal Q}_{1}$ differ in exactly one coordinate, then their associated codewords (where all the cycles are the same are also in the same shift) in the associated equivalence classes
@@
can be combined into one codeword by the merge-or-split method, i.e., they have a pair of companion vertices.
@@
These two codewords differs in two coordinates as is Lemma 14 and hence have a pair of companion states. ∎
We continue to consider two self-dual sequences of length $2^{t+1}$, ${\cal S}_{1}$ and ${\cal S}_{2}$ which differ in exactly two coordinates $i$ and $i+2^{t}$.
@@
Any two self-dual sequences of length $2^{t+1}$, ${\cal S}_{1}$ and ${\cal S}_{2}$ which differ in exactly two coordinates have in their associated codewords a pair of codewords whose vertices have a pair of companion vertices.
@@
The last important step is to use Lemmas 1415, and 16 to combine all the cyclic lists into one cyclic list which contains all the codewords and each two consecutive codewords have a pair of companion vertices.
@@
All the codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC $\mathbb{C}$ obtained in the construction of Theorem 9 can be ordered in a cyclic list in which each two consecutive codewords have a pair of companion vertices.
@@
We start by merging together two cyclic lists associated with the same self-dual sequence ${\cal S}$ and two different equivalence classes which are the first two in the list of equivalence classes associated with $Q_{2}$,
@@
and continue with the elements of the same equivalence class until it reaches
@@
The next element in the combined list will be
@@
which differ in exactly one coordinate from the first element
@@
We can continue in the same manner with the third equivalence class in the list in its usual order and the fourth equivalence class in reverse order and so on until all the equivalence classes associated with ${\cal S}$ are merged together.
Once all the codewords of the different equivalence classes for each self-dual sequence were merged together into one cyclic list in which each two consecutive codewords have a pair of companion vertices we have to combine the different codewords of distinct self-dual sequences into one long cyclic lict in which each two consecutive codewords have a pair of companion vertices. Let $S_{1}$ and $S_{2}$ be the first two self-dual sequences in the list. By the definition, ${\cal S}_{1}$ and ${\cal S}_{2}$ differ in exactly two coordinates $\ell$ and $\ell+2^{t}$. We take for these two self-dual sequences two equivalence classes that differ only in the coordinates in which ${\cal S}_{1}$ and ${\cal S}_{2}$ differ and merge the two list using the same idea as the one in merging the lists of two different equivalence classed for the same self-dual sequence.
Thus, the outcome is a cyclic list with all the codewords of $\mathbb{C}$, in which each two consecutive codewords have a pair of companion vertices. ∎
@@
If $k$ and $n$ are integers such that $k<n<2^{k}$, then there exist a $(2^{k},2^{t+1+s};n,2^{t})$-DBAC, where $10\leq s\leq n2^{t}-k-t-$ 1, of size $2^{n2^{t}-k-t-1-s}$.
An example for the list of codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC such that each two codewords have a pair of companion state is too large and hence we omit such an example and leave it for the interested reader.
A second example of the method is to start with the codewords of a $(2^{k},2^{t};n,2^{t}-1)$-DBAC $\mathbb{C}$ obtained from the construction of Theorem 8 with ${\cal F}$ as the PF$(n,k)$. This code has $2^{nm-k-t}$ codewords, where $m=2^{t}-1$ and $t\geq k$. We start with one partition of these codewords into equivalence classes (disjoint sets) as follows. Each codeword ${\cal C}\in\mathbb{C}$ is represented as a word of length $2^{t}$ as follows:
@@
where this representation is the one among the $2^{t}$ cyclic shifts of ${\cal C}$ with the smallest lexicographic value for $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ (although this lexicographic value will have no influence on the construction).
We define an equivalence relation on all the codewords of $\mathbb{C}$. The codeword ${\cal C}$ will be in the equivalence class
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}}),$
Each $X_{i_{j}}$, except for the last one, can be chosen in $2^{n-k}$ distinct ways for a total of $2^{(n-k)(2^{t}-1)}$ distinct choices. the last one is determined by the equation $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$. Since each one of the $2^{t}$ cycles from ${\cal F}$ in the equivalence class
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$
can be chosen as the first one, it follows that the number of equivalence classes in this equivalence relation is
@@
Merging the codewords in each equivalence class can be made exactly as in the first example. The first sequence is taken in its zero shift, The next $22^{t}-$ 2 sequences can be taken each one in any of its $2^{k}$ distinct cyclic shifts. The shift of the last sequence it determined in a way that the sum of all the shifts is congruent to 0 modulo $2^{k}$. Two codewords of an equivalence class which differ in exactly two consecutive shifts have a pair of companion states since the number of columns of a state is $12^{t}-$ 1 and these two codewords can be horizontally shifted in a way that they share the first $22^{t}-$ 2 columns and their shifts. For this purpose we consider the set $Q_{3}$ whose elements define the shifts of the sequences from ${\cal F}$. The set $Q_{3}$ is defined as follows.
@@
where the element $(0,j_{2},j_{3},\ldots,j_{2^{t}})\in Q_{3}$ for the equivalence class $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ is associated with the codeword
@@
Now, we should order the elements of $Q_{3}$ in a cyclic list such that each two elements in the list differ in exactly two consecutive elements. Such a list implies that each number of consecutive elements in the least can be merged together via the merge-or-split method. To have a longer cyclic list which includes all the codewords of $\mathbb{C}$ we have to order the equivalence classes in another cyclic list, where each two consecutive equivalence classes in the list differ in exactly two consecutive elements. Note, that these equivalence classes are over an alphabet of size $2^{n-k}$, the number of sequences in ${\cal F}$. Two codewords from consecutive elements in such a list which are associated with the same shifts from $Q_{3}$ differ in exactly two coordinates and hence have a pair of companion states. We do not have a general construction for such a list and we leave it as a problem for future research. When the number of sequence in ${\cal F}$ is two, i.e., $n-k=1$, it is relatively easy to generate such a list. The list consists of all the cyclic sequences of length $2^{t}$ and odd weight. We should order these elements in a list in which each two consecutive elements differ in exactly two coordinates. For this purpose we consider the list of self-dual sequences of length $2^{t+1}$. If $Y_{i}$ and $Y_{i+1}$ are two consecutive elements in this list, then $Y_{i}$ and $Y_{i+1}$ differ in two positions separated by $12^{t}-$ 1 other positions. ${\bf D}Y_{i}$ and ${\bf D}Y_{i+1}$ have odd weight and they differ in exactly two adjacent positions. Thus, the existence of a list for the self-dual sequences of length $2^{t+1}$ implies the list of sequences of odd weight and length $2^{t}$.
@@
If $n-k=1$, then there are two sequences in a PF$(n,k)$, $0001$ and $0111$ and these two sequences are the columns of all the codewords in $\mathbb{C}$. When $t=2$ the equivalence classes are also $0001$ and $0111$. The ordered list of these two equivalence classes is trivial. A list with all the elements of $Q_{3}$ is given as follows (the elements of $Q_{3}$ are presented as column vectors).
@@
The code $\mathbb{C}$ has 32 codewords. The 16 codewords of the equivalence class $0001$ are ordered in the list as follows.
@@
The 16 codewords of the equivalence class $0111$ are ordered in the list as follows.
@@
Now, it is easy to verify that any set of consecutive codewords can be merged together. Hence, we can obtain a $(4,2^{2+i};3,3)$-DBAC for each $i$, $0\leq i\leq 5$,
VI Conclusion

### v2

@@ V Joining Cycles of $G_{{\cal F},m}$
Our next step in combining the codewords is to order the $2^{k(2^{t}-1)}$ elements of ${\cal Q}_{1}$ in a cyclic Gray code, i.e., so that each two elements will differ in exactly one position. Such a cyclic Gray code is well-known [39] and is also very easy to construct. The following lemma can be readily verified.
@@
can be combined into one codeword using the merge-or-split method, i.e., they have a pair of companion vertices.
@@
Since the two codewords in $Q_{1}$ differ in exactly one coordinate, the associated two codewords of $\mathbb{C}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, differ in exactly two coordinates separated by $12^{t}-$ 1 coordinates in which both ${\cal C}_{1}$ and ${\cal C}_{2}$ are identical. The two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ associated with these $12^{t}-$ 1 coordinates and the coordinate which follows them differ only in the shift of the last column. This implies that these two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ are companion vertices. ∎
The existence of a cyclic Gray code of the elements of the set ${\cal Q}_{1}$ and Lemma 12 mean that we can form a cyclic Gray code from the codewords of each equivalence class.
The next step is to order the codewords from different equivalence classes associated with the same self-dual sequence ${\cal S}$. For this purpose we define another Gray code on the set of elements
@@
It is again easy to verify that we can find a pair of companion vertices in two distinct equivalence classes associated with the same self-dual sequence ${\cal S}$, which differ in exactly one coordinate of $Q_{2}$.
@@
If $(i_{i},i_{2},i_{3},\ldots,i_{2^{t}})\in{\cal Q}_{2}$ and $(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3},\ldots,i^{\prime}_{2^{t}})\in{\cal Q}_{1}$ differ in exactly one coordinate, then their associated codewords (where all the cycles are the same and are also in the same shift) in the associated equivalence classes
@@
can be combined into one codeword using the merge-or-split method, i.e., they have a pair of companion vertices.
@@
These two codewords differs in two coordinates as in Lemma 12, and hence have a pair of companion vertices. ∎
We next consider two self-dual sequences of length $2^{t+1}$, ${\cal S}_{1}$ and ${\cal S}_{2}$ which differ in exactly two coordinates $i$ and $i+2^{t}$.
@@
Any two self-dual sequences of length $2^{t+1}$, which differ in exactly two coordinates have in their associated codewords a pair of codewords whose vertices have a pair of companion vertices.
@@
The last step is to use Lemmas 1213, and 14 to combine all the cyclic lists into one cyclic list which contains all the codewords, where each two consecutive codewords have a pair of companion vertices.
@@
All the codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC $\mathbb{C}$ obtained in Construction 2 can be ordered in a cyclic list in which each two consecutive codewords have a pair of companion vertices.
@@
We start by merging together two cyclic lists associated with the same self-dual sequence ${\cal S}$ and two different equivalence classes, which are the first two in the list of equivalence classes associated with $Q_{2}$,
@@
and continues with the elements of the same equivalence class until it reaches
@@
The next element in the combined list is
@@
which differs in exactly one coordinate from the first element
@@
We continue in the same manner with the third equivalence class in the list in its usual order and the fourth equivalence class in reverse order and so on until all the equivalence classes associated with ${\cal S}$ are combined.
Once all the codewords of the equivalence classes for each self-dual sequence have been merged into one cyclic list in which each two consecutive codewords have a pair of companion vertices, we need to combine the different codewords of distinct self-dual sequences into one long cyclic list in which each two consecutive codewords have a pair of companion vertices. Let ${\cal S}_{1}$ and ${\cal S}_{2}$ be the first two self-dual sequences in the list. By definition, ${\cal S}_{1}$ and ${\cal S}_{2}$ differ in exactly two coordinates $\ell$ and $\ell+2^{t}$. For these two self-dual sequences, we take two equivalence classes that differ only in the coordinates in which ${\cal S}_{1}$ and ${\cal S}_{2}$ differ and merge the two lists using the same approach as the one used in merging the lists of two different equivalence classes for the same self-dual sequence.
The outcome is a cyclic list including all the codewords of $\mathbb{C}$, in which each two consecutive codewords have a pair of companion vertices. ∎
@@
If $t$, $k$, $\ell$, and $n$ are integers such that $k<n<2^{k}$, $1\leq\ell$, and $\ell+1\leq t\leq n2^{\ell}-k$, then there exists a $(2^{k},2^{t};n,2^{\ell})$-DBAC of size $2^{n2^{\ell}-k-t}$.
A nontrivial example ($t>1$) for the list of codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC with the property that each two codewords have a pair of companion vertices is too large to include here and hence we omit such an example and leave it for the interested reader.
A second example for the idea of joining cycles in $G_{{\cal F},m}$ starts with the codewords of a $(2^{k},2^{t};n,2^{t}-1)$-DBAC $\mathbb{C}$ obtained from Construction 1 with ${\cal F}$ as the PF$(n,k)$ which are ordered arbitrarily by $X_{0},X_{1},\ldots,X_{2^{n-k}-1}$. Such a code has $2^{nm-k-t}$ codewords, where $m=2^{t}-1$ and $t\geq k$. Partition these codewords into equivalence classes (disjoint sets) as follows. Each codeword ${\cal C}\in\mathbb{C}$ is represented as a word of length $2^{t}$ as follows:
@@
where this representation is the one among the $2^{t}$ cyclic shifts of ${\cal C}$ with the smallest lexicographic value for $(i_{1},i_{2},\ldots,i_{2^{t}})$ (although this lexicographic value has no influence on the construction). The codeword ${\cal C}$ is in the equivalence class
$(i_{1},i_{2},\ldots,i_{2^{t}}),$
Each $X_{i_{j}}$, except for the last, can be chosen in $2^{n-k}$ distinct ways for a total of $2^{(n-k)(2^{t}-1)}$ distinct choices. The last is determined by the equation $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$. Since each of the $2^{t}$ cycles from ${\cal F}$ in the equivalence class
$(i_{1},i_{2},\ldots,i_{2^{t}})$
can be chosen as the first, it follows that the number of equivalence classes in this equivalence relation is
@@
Merging the codewords in each equivalence class can be achieved exactly as in the first example. The first sequence is taken in its zero shift. The next $22^{t}-$ 2 sequences taken in any of the $2^{k}$ distinct cyclic shifts. The shift of the last sequence is chosen so that the sum of all the shifts is congruent to 0 modulo $2^{k}$. Two codewords of an equivalence class which differ in exactly two consecutive shifts have a pair of companion vertices since the number of columns of the matrix associated with the vertex is $12^{t}-$ 1 and these two codewords can be horizontally shifted in a way that they share the first $22^{t}-$ 2 columns and their shifts. For this purpose we consider the set $Q_{3}$, whose elements define the shifts of the sequences from ${\cal F}$, is defined as follows.
@@
where the element $(0,j_{2},j_{3},\ldots,j_{2^{t}})\in Q_{3}$ for the equivalence class $(i_{1},i_{2},\ldots,i_{2^{t}})$ is associated with the codeword
@@
Order the elements of $Q_{3}$ in a cyclic list such that each pair of consecutive elements in the list differ in exactly two consecutive elements. Such a list implies that each number of consecutive elements in the list can be combined via the merge-or-split method. To have a longer cyclic list which includes all the codewords of $\mathbb{C}$ we need to order the equivalence classes in another cyclic list, where each two consecutive equivalence classes in the list differ in exactly two consecutive elements. Note, that these equivalence classes are over an alphabet of size $2^{n-k}$, the number of sequences in ${\cal F}$. Two codewords from consecutive elements in such a list which are associated with the same shifts from $Q_{3}$ differ in exactly two coordinates and hence have a pair of companion vertices.
We do not have a general construction for such a list and we leave it as a problem for future research. However, when the number of sequences in ${\cal F}$ is two, i.e., $n-k=1$, it is relatively easy to generate such a list. The list consists of all the cyclic sequences of length $2^{t}$ and odd weight. Order these elements so that each two consecutive elements differ in exactly two coordinates. For this purpose we consider the list of self-dual sequences of length $2^{t+1}$. If $Y_{i}$ and $Y_{i+1}$ are two consecutive elements in this list, then $Y_{i}$ and $Y_{i+1}$ differ in two positions separated by $12^{t}-$ 1 other positions. ${\bf D}Y_{i}$ and ${\bf D}Y_{i+1}$ have odd weight and they differ in exactly two adjacent positions. Thus, the existence of a list for the self-dual sequences of length $2^{t+1}$ implies the list of sequences of odd weight and length $2^{t}$.
@@
If $n-k=1$, then there are two sequences in a PF$(n,k)$: $[0001]$ and $[0111]$, and these two sequences are the columns of all the codewords in $\mathbb{C}$. When $t=2$ the equivalence classes are $(0,0,0,1)$ and $(0,1,1,1)$. The ordered list of these two equivalence classes is trivial. A list with all the elements of $Q_{3}$ is given as follows (the elements of $Q_{3}$ are presented as column vectors).
@@
The code $\mathbb{C}$ has 32 codewords. The 16 codewords of the equivalence class $(0,0,0,1)$ are ordered in the list as follows.
@@
The 16 codewords of the equivalence class $(0,1,1,1)$ are ordered in the list as follows.
@@
It is easy to verify that any set of consecutive codewords can be combined. Hence, we obtain a $(4,2^{2+i};3,3)$-DBAC for each $i$, $0\leq i\leq 5$.
$\blacksquare$
VI Analysis of the Constructions
In this section we summarize the parameters of DBACs which are obtained by our direct and recursive constructions including the joining of cycles via the merge-or-split method. We further apply the recursive constructions to characterize some families of DBACs which were completely solved by our constructions.
We start with the existence of perfect factors which play an important role in our constructions. Their weight of their sequences is also have an important role in the constructions. We can analyze their existence with either sequences of odd weight or even weight. For sequences of even weight we can characterize all the parameters of perfect factors.
Theorem 14
  1. 1.
  2. 2.
Proof:
  1. 1.
    If $k\leq n<2^{k-1}$, then consider a span $k$ de Bruijn sequence ${\cal S}$ with linear complexity $2^{k-1}+k$ whose existence is implied by Theorem 1. By Lemmas 79, and 10 we have that ${\bf D}^{-r}{\cal S}$ for any $r$ such that $0\leq r<2^{k-1}-k$ contains $2^{r}$ sequences of length $2^{k}$ in $G_{k+r}$, i.e., $2^{n-k}$ sequences of length $2^{k}$ in $G_{n}$ which is a PF$(n,k)$ for any given $n$ such that $k\leq n<2^{k-1}$. By Lemma 7 all the generated sequences have linear complexity $2^{k-1}+k+r<2^{k}$ and hence by Lemma 5 all these sequences have even weight.
  2. 2.
    For $n=2^{k}-1$ let ${\cal F}$ be a perfect factor PF$(n,k)$ with sequences of length $n+1=2^{k}$. One of these sequences contains the all-zero word of length $n$. Clearly, the sequence which contains the all-zero word of length $n+1$ is $[0^{n}1]$. This sequence contains all words of length $n$ and weight one. Since each cyclic sequence of length $n+1$ and weight two contains exactly two words of weight one (by omitting a one from the sequence), it follows that ${\cal F}$ does not contain sequences of weight two and hence all words of weight two are contained in sequences with weight three. This implies that ${\cal F}$ does not contain sequences of weight four and so on by induction. Thus, there is a unique factor whose sequences have odd weight. This factor is the state diagram of the CSRn.
Theorem 14 implies that whenever a perfect factor PF$(n,k)$ exists, except for $n=2^{k}-1$, i.e., when $k\leq n<2^{k}-1$ by Theorem 2, there exists a perfect factor for which all its sequences have even weight. The situation is more complicated for odd weight. PF$(n,n)$ contains exactly one sequence of length $2^{n}$ in $G_{n}$, i.e., a span $n$ de Bruijn sequence. Hence, its weight is always even. For all the other parameters a perfect factor in which all the sequences have odd weight could potentially exist. Unfortunately, whether such perfect factors exist is an open problem. By Theorem 14, if $n=2^{k}-1$, then all the sequences in a PF$(n,k)$ have odd weight. For the other parameters we can use the same technique used in the proof of Theorem 14; the details are left for the reader. In other words, we can start with a span $k$ de Bruijn sequence whose linear complexity is $2^{k-1}+\delta$, where $k\leq\delta\leq 2^{k-1}-k-1$ and $\delta\neq k+1$, and apply the inverse operator ${\bf D}^{-r}$, $r>0$, until we reach sequences with linear complexity $2^{n}$, where $n\geq k$. The sequences obtained will be of odd weight by Lemma 5. Other than these two techniques we do not know how to construct perfect factors whose sequences all have odd weights.
It is important to note that in all the constructions mentioned so far, if a sequence ${\cal S}$ is contained in the perfect factor ${\cal F}$, then $\bar{{\cal S}}$ is also a sequence in ${\cal F}$ (this includes self-dual sequences), unless ${\cal F}$ contains exactly one sequence. By Lemma 6 all these sequences of length $2^{k}$ are self-dual if and only if they are of linear complexity $12^{k-1}+$ 1. However, perfect factors without these properties can also be constructed.
Example 2
$\blacksquare$
The following result is a consequence from the discussion in Section II-D, and implies that the requirement for a perfect factor in Construction 2 can be always satisfied.
Lemma 15
[…]
Proof:
The methods used to generate perfect factors in Theorem 14 also allows us to generate perfect factors without self-dual sequences, and if ${\cal F}$ is the PF$(n,k)$, then ${\cal S}\in{\cal F}$ if and only if $\bar{{\cal S}}\in{\cal F}$. This is apart for the case when $n=2^{k-1}$ and the sequences of PF$(n,k)$ have linear complexity $12^{k-1}+$ 1. In this case by Lemma 6 the obtained sequences are self-dual.
[…]
We continue by summarizing the parameters of the constructions which were presented. For the direct constructions the parameters are presented in Table I. For the recursive constructions the parameters are presented in Table II.
construction parameters perfect factor constructed DBAC size
Construction 1 $m=2^{t}-1$ PF$(n,k)$ $(2^{k},2^{t};n,m)$ $2^{nm-k-t}$
Corollary 4 $1\leq\ell$, $k<n<2^{k}$, $\ell+1\leq t\leq n2^{\ell}-k$ PF$(n,k)$ no self-dual $(2^{k},2^{t};n,2^{\ell})$ $2^{n2^{\ell}-k-t}$
Construction 3 $1\leq t\leq n-k$ PF$(n,k)$ $(2^{k},2^{t};n,1)$ $2^{n-k-t}$
[…]
TABLE I: Direct constructions of DBACs
construction columns’ weight parameters perfect factor constructed DBAC size
Construction 6 even $2\leq k,n,m$, $t=m-\ell$ PF$(m,m-\ell)$ $(2^{k},2^{m-\ell};n+1,m)$ $2^{m}\Delta$
$0\leq\ell$, $m<2^{m-\ell}$
Construction 7 even $2\leq k,n,t,m$ PF$(m,m)$ $(2^{k},2^{m+t};n+1,m)$ $\Delta$
Construction 8 even $3\leq m<2^{t}$, $2\leq k,n$ PF$(m,m-1)$ $(2^{k},2^{m+t-1};n+1,m)$ $2\Delta$
Construction 9 odd $3\leq m<2^{t}$, $2\leq n<2^{k}$ PF$(m,t)$ no self-dual $(2^{k+1},2^{m};n+1,m)$ $2^{m-1}\Delta$
Construction 10 odd $3\leq m<2^{t}$, $2\leq n<2^{k}$ PF$(m,m-1)$ $(2^{k+1},2^{m+t-1};n+1,m)$ $\Delta$
TABLE II: Recursive constructions with $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$
What new parameters of DBACs can be obtained by the recursive construction? For example, what DBACs with new parameters can be obtained using Construction 5? Direct constructions can be used to form $(2^{k},2^{t};n,m)$-DBACs, where the columns of the codewords are sequences from a PF$(n,k)$ say ${\cal F}$. Theorem 7 implies the existence of a $(2^{k},2^{t};n+1,m)$-DBAC. But, such a DBAC can be obtained using a direct construction with a PF$(n+1,k)$ that can be obtained from ${\bf D}^{-1}{\cal F}$. So applying the recursion in this way does not yield DBACs with new parameters. On the other hand we can transpose the DBACs obtained from the direct constructions and use the recursion to obtained DBACs with new parameters. For this purpose, we have to examine the weight of the rows in the direct constructions. To demonstrate the necessary analysis we concentrate on Construction 2.
Construction 2 is used to obtain $(2^{k},2^{t+1};n,2^{t})$-DBACs. The codewords are $2^{k}\times 2^{t+1}$ matrices, where the $i$-th column is the complement of the $(2^{t}+i)$-th column. In other words, each row is a self-dual sequence whose linear complexity is $12^{t}+$ 1, which implies that its weight is even. Moreover, applying ${\bf D}^{-\ell}$ on the sequences up to $\ell\leq 2^{t}-2$ yields sequences of even weight, so we will obtain DBACs with new parameters using Construction 2. Applying ${\bf D}^{-2^{t}+1}$ yields sequences with linear complexity $2^{t+1}$ which are sequences of odd weight. This makes this construction very important for the recursive constructions.
Specifically, Let $\mathbb{C}$ be a $(2^{k},2^{t+1};n,2^{\ell})$-DBAC obtained by joining the codewords of Construction 2. A codeword obtained in Construction 2 is a $2^{k}\times 2^{t+1}$ doubly-periodic array of the form $[X~{}\bar{X}]$, where $X$ is a $2^{k}\times 2^{t}$ array. Therefore, each row of the array is a self-dual sequence whose linear complexity is $12^{t}+$ 1. Combining two such codewords yields an array of the form $[X~{}\bar{X}~{}Y~{}\bar{Y}]$. Each row of such array is of length $2^{t+2}$ and its linear complexity is at most $2^{t+1}+2^{t}$. If we continue to combine codewords for a $(2^{k},2^{s};n,2^{t})$-DBAC, where the linear complexity of a row in the obtained array is at most $2^{s}-2^{t}$. The codewords obtained can be transposed to obtain a $(2^{s},2^{k};2^{t},n)$-DBAC whose columns have even weight and linear complexity at most $2^{s}-2^{t}$. On this code we can apply $j$ iterations of Construction 7, where $11\leq j\leq 2^{t}-$ 1 to obtain a $(2^{s},2^{k};2^{t}+j,n)$-DBAC. The obtained codewords of the code can be transposed again to obtain a new $(2^{k},2^{s};n,2^{t}+j)$-DBAC $\mathbb{C}^{\prime}$. The codeword of $\mathbb{C}^{\prime}$ can be merged by using the merge-or-split method to obtain more DBACs with larger codewords and a smaller size.
Similar combinations of the direct and recursive constructions with a combination of the merge-or-split method as described in Section V yield more families of DBACs with various parameters. A few families of $(2^{k},2^{t};n,m)$-DBACs obtained from the constructions are summarized in table III.
DBAC parameters constructions/reference
$(2^{k},2^{t};n,1)$ $k\leq n<2^{k}$, $1\leq t\leq n-k$ Construction 3
$(2^{k},2^{t};n,2^{\ell})$ $1\leq\ell$, $k\leq n<2^{k}$, $\ell+1\leq t\leq n2^{\ell}-k$ joining cycles/Corollary 4
$(2^{k},2^{s};n,2^{t}+j)$ $1\leq t$, $10\leq j\leq 2^{t}-$ 1, $k\leq n<2^{k}$, $t+1\leq s\leq n2^{t}-k$ joining cycles/Corollary 4/Construction 6
TABLE III: Infinite families of DBACs
VII Conclusion and Future Work